OpenStudy (mackenzie6302):

Which is equal to –(4 – 5)3 + 1? 0 2 1 –1

alones (alones):

@Mackenzie6302 −(4−5))(3)+1 =(−(−1))(3)+1 =(1)(3)+1 =3+1 =4 hoep steps help

Directrix (directrix):

Is the posted expression meant to be the following: –(4 – 5)^3 + 1 If you crank that out correctly, you will get one of the answer options. @Mackenzie6302

alones (alones):

@Directrix but umm did i do it right??? /:

Directrix (directrix):

@AloneS −(4−5)*(3)+1= 4, yes. It is not one of the posted answer options. I think that the poster meant to write –(4 – 5)³ + 1 which does not equal 4.

alones (alones):

YAASS!!! Yea I think so too !!!!! At least I got somethign right form yesturday!!